Find the following probabilities:

a) Pr{Z < 0.66}

b) Pr{Z ≥ -0.66}

c) Pr{-2.01 < Z < 2.01}

d) Pr{-1.91 < Z < 0.0}

e) Pr{Z < -1.35 or Z > 1.35} (you want the probability that Z is outside the range -1.03 to 1.03)
Answer #1

a) P(Z<0.66) = (0.66) 0.745373

b) P(Z\geq -0.66) = 1 - P(Z<-0.66)=1- \Phi(-0.66) = 1 - 0.254627 = 0.745373

c)

P(-2.01 Z <2.01) P(Z < 2.01) P(Z <-2.01)

\Rightarrow P(-2.01<Z<2.01) = \Phi(2.01) - \Phi(-2.01)

P-2.01 Z 2.01)=1 - -2.01)-(2.01)

\Rightarrow P(-2.01<Z<2.01) = (1 - 0.0222156) - 0.0222156

\Rightarrow P(-2.01<Z<2.01) = 0.955569

d)

P-1.91 Z 0) P(Z <0) P(Z<-1.91) (0)-P(1.91)

\Rightarrow P(-1.91 <Z<0) = 0.5 - 0.0280666

\Rightarrow P(-1.91 <Z<0) = 0.471933

e)

P(Z<-1.35,Z>1.35) = P(Z<-1.35) + P(Z>1.35)

\Rightarrow P(Z<-1.35,Z>1.35) = 2*P(Z<-1.35)

\Rightarrow P(Z<-1.35,Z>1.35) = 2*0.0885080

\Rightarrow P(Z<-1.35,Z>1.35) = 0.177016